a constant volume gas thermometer exerts an absolute pressure of 1000Pa at room temperature. if the pressure has been increased to 1500Pa. what will be the new temperature?


1
Expert's answer
2021-10-18T11:11:03-0400

The Charles's law (constant volume processes) says

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

Hence, the final temperature

"T_2=\frac{P_2}{P_1}T_1=\frac{1500\:\rm Pa}{1000\:\rm Pa}(20+273)=440\:\rm K=166^{\circ}C"